OpenStudy (anonymous):

Becky and Tommy are asked to solve −4x − 25 = 8x + 15. Identify where one of them made an error. Becky made the error when she subtracted 8x Becky made the error when she divided by 40 Tommy made the error when he added 4x Tommy made the error when he added 15

OpenStudy (anonymous):

Becky −4x − 25 = 8x + 15 −4x − 25 − 8x = 8x + 15 − 8x −12x − 25 = 15 -12x − 25 + 25 = 15 + 25 −12x = 40 Negative 12x over 40 equals 40 over 40 x=negative three-tenths

OpenStudy (anonymous):

tommy −4x − 25 = 8x + 15 −4x − 25 + 4x = 8x + 15 + 4x −25 = 12x + 15 −25 − 15 = 12x + 15 − 15 −40 = 12x negative 40 over 12 equals 12x over 12 x=negative ten-thirds

OpenStudy (anonymous):

please help @princeharryyy

OpenStudy (princeharryyy):

checking ur solution ... @kimberlylopez

OpenStudy (princeharryyy):

why do u think tommy did it

OpenStudy (princeharryyy):

@kimberlylopez

OpenStudy (princeharryyy):

u there @kimberlylopez

OpenStudy (anonymous):

yeahh ive got the answer thanks anyway

OpenStudy (princeharryyy):

whats ur answer @kimberlylopez

OpenStudy (anonymous):

Becky made the error when she divided by 40

OpenStudy (princeharryyy):

how ????

OpenStudy (anonymous):

because your trying to get anyway for the x ?.

OpenStudy (princeharryyy):

have they given this ... Becky −4x − 25 = 8x + 15 −4x − 25 − 8x = 8x + 15 − 8x −12x − 25 = 15 -12x − 25 + 25 = 15 + 25 −12x = 40 Negative 12x over 40 equals 40 over 40 x=negative three-tenths tommy −4x − 25 = 8x + 15 −4x − 25 + 4x = 8x + 15 + 4x −25 = 12x + 15 −25 − 15 = 12x + 15 − 15 −40 = 12x negative 40 over 12 equals 12x over 12 x=negative ten-thirds

OpenStudy (princeharryyy):

@kimberlylopez

OpenStudy (princeharryyy):

u dere @kimberlylopez

OpenStudy (princeharryyy):

???? be quick cos ur answer may be wrong i am suspecting it @kimberlylopez

OpenStudy (princeharryyy):

now if you are given those equations as above the answer is gonna be tommy who did it rather then sweet becky @kimberlylopez

OpenStudy (princeharryyy):

and the answer is gonna be D @kimberlylopez